Find the dual program of the following linear programming problem.

π‘€π‘–π‘›π‘–π‘šπ‘–π‘§π‘’ 𝑍 = 30π‘₯1 βˆ’ 50π‘₯2 + 10π‘₯3

𝑠𝑒𝑏𝑗𝑒𝑐𝑑 π‘‘π‘œ

3π‘₯1 + 2π‘₯2 βˆ’ π‘₯3 β‰₯ 44

π‘₯1 βˆ’ π‘₯2 + π‘₯3 = 7

π‘₯1 𝑖𝑠 π‘’π‘›π‘π‘œπ‘›π‘ π‘‘π‘Ÿπ‘Žπ‘–π‘›π‘’π‘‘, π‘₯2 β‰₯ 0, π‘₯3 β‰₯ 0,


1
Expert's answer
2021-09-27T14:14:45-0400

In order to apply the dual simplex method, convert Min Z to Max Z and allΒ β‰₯Β constraint to ≀ constraint by multiply -1.


MaxΒ Z=-30x1+50x2-10x3

subject to

-3x1-2x2+x3≀-44

x1-x2+x3=7

and Β x1 𝑖𝑠 π‘’π‘›π‘π‘œπ‘›π‘ π‘‘π‘Ÿπ‘Žπ‘–π‘›π‘’π‘‘,Β x2,x3β‰₯0;


The problem is converted to canonical form by adding slack, surplus and artificial variables as appropiate


1. As the constraint-1 is of type '≀' we should add slack variableΒ S1


2. As the constraint-2 is of type '=' we should add artificial variableΒ A1


After introducing slack,artificial variables

MaxΒ Z=-30x1+50x2-10x3+0S1-MA1

subject to

-3x1-2x2+x3+S1=-44

x1-x2+x3+A1=7

andΒ x1,x2,x3,S1,A1β‰₯0


Here not allΒ Zj-Cjβ‰₯0.Β (BecauseZ1-C1=-M+30)

Hence, method fails to get optimal basic feasible solution.